OpenStudy (anonymous):

I need help with Equations and Slopes. The directions are to graph each line using the slope and the y intercept. the problem is: 2y - 6x= -8

OpenStudy (anonymous):

first put the equation into y=mx+b form (solve for y)

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

you should get y=3x+4. is that what you got?

OpenStudy (anonymous):

How did you get the 3 and the 4?

OpenStudy (campbell_st):

pt slope form y = 3x + 4... let x = 0 to get then solve for y to get the y- intercept... Next let y = 0, solve for x to get the x - intercept

OpenStudy (anonymous):

get the y by itself: 2y-6x=8. First add 6x to both sides. 2y=6x+8 then divide by 2 to get the y alone: y=3x+4

OpenStudy (anonymous):

oh okay, what next

OpenStudy (anonymous):

y=mx+b: y=3x+4. b= the y-intercept (where the line will cross the y axis) and the slope is m. m=3 (remember that its 3/1) up 3, over 1. or down 3 over 1

OpenStudy (anonymous):

Thank you so much, if I have any further questions in the future i'll be sure to message you.

OpenStudy (anonymous):

no problem